Silver (107.9 g/mol) is electroplated onto a key using a current of 2.00 amp for a period of 30.0 minutes. What mass of silver will be deposited onto the key?

Ag+(aq) + e = Ag(s) E° = +0.80 V
Answer #1

Given molar mass of Ag =107.9g/mol

Ag++ e- -----> Ag(107.9g)

From the equation it is clear that to coat 107.9g of Ag

we need one mole of electron. Charge of one mole of electron=96500 Coulombs .Charge of one mole of electron is known as farad(F)

1F =96500 coulombs

Current used =2 Amps (given)

Time period for which current is applied=30 minutes= 30×60=1800sec

Total charge = current used ×time period of use of current

So total charge =2×1800sec =3600 Coulomb

Mass of Ag coated = molar mass of Ag ×(total charge/96500)

=107.9g×(3600C/96500C) =4.02g

Mass of silver coated=4.02g

Thank you